# UniTS: Unified Time Series Generative Model for Remote Sensing

### v1

@@ Abstract
One of the primary objectives of satellite remote sensing is to capture the complex dynamics of the Earth environment, which encompasses tasks such as reconstructing continuous cloud-free time series images, detecting land cover changes, and forecasting future surface evolution. However, existing methods typically require specialized models tailored to different tasks, lacking unified modeling of spatiotemporal features across multiple time series tasks. In this paper, we propose a Unified Time Series Generative Model (UniTS), a general framework applicable to various time series tasks, including time series reconstruction, time series cloud removal, time series semantic change detection, and time series forecasting. Based on the flow matching generative paradigm, UniTS constructs a deterministic evolution path from noise to targets under the guidance of task-specific conditions, achieving unified modeling of spatiotemporal representations for multiple tasks. The UniTS architecture consists of a diffusion transformer with spatio-temporal blocks, where we design an Adaptive Condition Injector (ACor) to enhance the model’s conditional perception of multimodal inputs, enabling high-quality controllable generation. Additionally, we design a Spatiotemporal-aware Modulator (STM) to improve the ability of spatio-temporal blocks to capture complex spatiotemporal dependencies. Furthermore, we construct two high-quality multimodal time series datasets, TS-S12 and TS-S12CR, filling the gap of benchmark datasets for time series cloud removal and forecasting tasks. Extensive experiments demonstrate that UniTS exhibits exceptional generative and cognitive capabilities in both low-level and high-level time series tasks. It significantly outperforms existing methods, particularly when facing challenges such as severe cloud contamination, modality absence, and forecasting complex phenological variations. More details can be found on the project page: https://yuxiangzhang-bit.github.io/UniTS-website/.
@@ I Introduction
Satellite image time series have become an indispensable tool for understanding and monitoring the dynamics of Earth’s systems [1], as they provide continuous spatiotemporal observational data. These long-term and consistent records of land surface dynamics deliver critical support for numerous fields, including ecological environment assessment [2], climate mitigation [3], and emergency response to natural disasters [4]. With the increasing availability of high spatiotemporal resolution satellite data, time series analysis has expanded beyond single-sensor or single-modal approaches to integrate multispectral, radar, panchromatic, and other multi-source data, enabling more comprehensive and accurate perception of land surface changes.
@@
Low-level tasks such as gap filling and cloud removal are essential and critical steps in the analysis and application of optical remote sensing images. This necessity stems from the inherent susceptibility of optical satellite imagery to atmospheric conditions. Interferences such as clouds, haze, and cloud shadows can partially or completely block optical signals reflected from the Earth’s surface, leading to missing pixels or radiometric distortion in the data received by the sensor, thereby severely limiting data usability. According to statistics from long-term observations by the MODIS sensor [5], approximately 67% of the Earth’s surface and 55% of the land area are persistently affected by cloud cover on average. This high coverage rate not only significantly reduces the practical observation efficiency of optical satellites but also makes completely cloud-free time series images extremely scarce at regional scales. Beyond weather factors, issues such as transient sensor failures or data transmission errors during satellite operation can also lead to data quality problems like strip loss, abnormal noise, or local data corruption. The primary goal of time series reconstruction is to recover complete, high-quality image time series from incomplete or noisy observational data. In current research, researchers often use cloud detection methods or extract binary cloud masks from existing data products (e.g., the cloud masks generated by the s2cloudless algorithm for Sentinel-2) and apply them to gap-free time series to simulate varying degrees of data missingness. Patrick et al. [6] proposed UnCRtainTS, which integrates an attention architecture with a multivariate uncertainty prediction method. UnCRtainTS not only calibrates prediction uncertainty but also enables precise control over reconstruction quality. Stucker et al. [7], proposed U-TILISE from the perspective of representation learning, employing a phased framework with an encoder-decoder and temporal attention encoder. It implicitly captures spatiotemporal patterns of spectral intensity from a representation learning standpoint, efficiently mapping cloud-masked input sequences to cloud-free outputs while balancing spatial detail preservation and temporal correlation mining. From a conditional generation viewpoint, Shu et al. [8] proposed the multimodal diffusion framework RESTORE-Diffusion Transformer (DiT). It achieves sequence-level optical-Synthetic Aperture Radar (SAR) fusion via a diffusion framework, utilizing SAR time series temporally matched with Sentinel-2 to provide subsurface dynamics under clouds and guide time series reconstruction, while embedding date information to address irregular observation intervals and periodic variations.
@@
Time series cloud removal is closely related to time series reconstruction and poses greater challenges. Unlike simply reconstructing areas obscured by cloud masks, this task specifically aims at restoring pixels occluded by clouds and their shadows within a temporal sequence. Patrick et al. [9] constructed the multi-temporal and multimodal benchmark dataset SEN12MS-CR-TS, where each region of interest (ROI) consists of 30 temporally aligned Sentinel-1 and Sentinel-2 images from 2018, paired with corresponding spatiotemporal patches. Furthermore, they implemented a 3D encoder-decoder architecture to achieve translation from multi-temporal Sentinel-1 to multi-temporal Sentinel-2 data. To address data gaps in visible and near-infrared bands, Gonzalez-Calabuig et al. [10] proposed GANFilling based on Generative Adversarial Networks (GANs). This approach incorporates convolutional long short-term memory layers into the GAN framework, enabling the transformation of cloudy optical sequences into cloud-free optical sequences.
@@
The continuous spatiotemporal time series data used for understanding and analysis in high-level tasks is provided by low-level tasks. Time series semantic change detection is a typical high-level understanding task used to monitor complex land cover dynamics. Unlike traditional binary change detection [11, 12], it not only identifies the specific locations where land surface changes occur over time but also further determines the semantic class and transition type of each changed area. In other words, it determines not only the precise location (Where) and timing (When) of changes but also provide semantic information about the type of change (What). Garnot et al. [13] proposed a U-Net with Temporal Attention Encoder (U-TAE) for time series panoptic segmentation. However, this method is only designed to predict a single land cover map corresponding to the time series images and does not support land cover change detection. He et al. [14] introduced an end-to-end framework based on a fully convolutional network, which directly learns the mapping between spectral features and land cover classes to achieve time series semantic change detection in a unified manner. Nevertheless, this approach primarily relies on fully convolutional networks for feature extraction. By irreversibly compressing spatial dimensions, it overlooks the importance of spatial and spectral information, particularly neglecting explicit spatiotemporal modeling.
Time series forecasting involves analyzing historical data of the Earth’s surface and its related influencing factors to identify change patterns and predict future surface conditions. This task requires modeling complex temporal evolution patterns, including learning the mapping relationships from historical observations to future scenarios, as well as capturing seasonal trends, periodic patterns, and long-term spatiotemporal dependencies. It predicts multiple future images based on a given set of current images. Requena-Mesa et al. [15] are the first to define the Earth surface forecasting task and construct the EarthNet2021 dataset, which includes four bands of Sentinel-2 visible and near-infrared data at a 20-meter resolution, matching topography and mesoscale (1.28 km) meteorological variables. Currently, one of the most representative works is Pangu-Weather proposed by Bi et al. [16]. By employing a 3D deep network equipped with specific priors and a hierarchical temporal aggregation strategy, it effectively handles complex patterns in weather data and reduces cumulative errors. Trained on 39 years of global data, Pangu-Weather demonstrates deterministic forecasting capabilities and performs well in extreme weather and ensemble forecasting. Benson et al. [17] constructed the first high-resolution vegetation forecasting dataset, GreenEarthNet, and proposed a multimodal transformer model named Contextformer. This model utilizes spatial context through a visual backbone and leverages a temporal transformer to predict the temporal evolution of context patches.
@@
  • Limited exploration in time series cloud removal tasks and challenges in dataset construction. Most current low-level studies focus primarily on time series reconstruction tasks, where cloud masks are used to simulate missing data or cloud cover, aiming to reconstruct surface reflectance in the masked areas. However, real-world scenarios with varying degrees of cloud contamination are far more complex than the assumptions underlying cloud mask simulations. In contrast, existing studies have made limited progress on the more challenging task of time series cloud removal, and the currently available benchmark datasets exhibit significant shortcomings. For instance, in the SEN12MS-CR-TS dataset [9], temporal misalignment between Sentinel-1 and Sentinel-2 modalities can be as long as 14 days, while the temporal gap between input data and ground truth may extend up to one year, introducing significant noise. Moreover, the dataset excludes images with over 50% cloud coverage for cloud removal evaluation. Similarly, EarthNet2021 dataset [15] omits sequences with heavy cloud cover, failing to meet the requirements of realistic cloud removal scenarios. Consequently, the absence of high-quality, temporally aligned time series datasets severely hinders reliable model training and evaluation in this field.
@@
  • Limited exploration in time series forecasting tasks. Most existing studies employ discriminative models, such as ConvLSTM and 3D CNN, with Pangu-Weather being the most representative work [16]. However, changes in geospatial-temporal data involve complex spatiotemporal state evolution, and discriminative models optimized based on mean-absolute-error loss struggle to effectively capture spatiotemporal dynamic distributions. Generative models with strong capability to fit complex distributions need to be further explored and applied in time series forecasting research. Furthermore, existing work primarily focuses on predicting vegetation indices, while forecasting of original remote sensing images/reflectance remains relatively scarce, particularly for multispectral imagery like Sentinel-2 with high spatiotemporal resolution. Due to its high spectral dimensionality and complex spatial details, predicting original multispectral time series presents even greater challenges.
@@
  • The absence of a unified framework capable of handling multiple remote sensing time-series tasks. Current research remains largely in the stage of developing specialized models tailored for specific tasks, and a unified framework that can effectively address diverse remote sensing time series tasks is still missing. We argue that the core challenge for both low-level and high-level time series tasks lies in mining spatiotemporal representations, with the main difference being how these representations are adapted through task-specific designs to meet particular requirements.
@@
To address the aforementioned technical and data gaps, we propose Unified Time Series Generative Model (UniTS) based on the flow matching-based generative paradigm, which is applicable to four time series tasks, as shown in Fig.1. Additionally, we construct two high-quality benchmarks TS-S12 and TS-S12CR for time series model validation. The main contributions of this paper are as follows:
@@
  • UniTS achieves a unified modeling for the first time in four types of tasks: time series reconstruction, time series cloud removal, time series semantic change detection, and time series forecasting. The framework demonstrates excellent generation and understanding capabilities across different task levels, whether it is low-level time series reflectance recovery or high-level time series semantic understanding.
@@
  • Within the diffusion transformer framework, we introduce a spatio-temporal block and design two novel components: the Adaptive Condition Injector (ACor) and the Spatiotemporal-aware Modulator (STM). ACor adaptively injects multimodal conditional information (e.g., SAR and optical imagery) by dynamically generating affine transformation parameters, significantly enhancing the model’s conditional perception of multimodal inputs across various time series tasks. Meanwhile, STM modulates attention weights in the spatio-temporal block by leveraging generated dynamic bias terms based on spatiotemporal priors, thereby strengthening the model’s capacity to capture complex spatiotemporal dependencies.
@@
  • We construct two high-quality multimodal time-series datasets, namely TS-S12 and TS-S12CR. Among them, TS-S12 and TS-S12CR contain Sentinel-1 and Sentinel-2 imagery from 14,973 and 12,126 ROIs around the world, respectively. TS-S12 provides aligned sample pairs of Sentinel-1 and cloud-free Sentinel-2 for time series reconstruction and forecasting tasks. TS-S12CR offers aligned triplet samples of Sentinel-1, cloud-covered Sentinel-2, and cloud-free Sentinel-2 specifically designed for time series cloud removal task. TS-S12CR provides an extreme scenario with an average cloud coverage of 84.02%, serving as an important benchmark for developing robust time series cloud removal methods.
@@ III Unified Time Series Generative Model
In this section, we first introduce the framework overview of UniTS in Section III-A. The details of ACor and STM are presented in Section III-D and III-E, respectively. Finally, we elaborate on the training and inference processes for conditional time series generation in Section III-E. Notations used in this paper are summarized in Table II.
@@ III-C UniTS Framework
Architecture. We instantiate ${f_{\theta}}(\cdot)$ in Eq.(2) using a transformer-based architecture. This architecture is chosen for its simplicity, scalability, and effectiveness in generative modeling. Specifically, UniTS is implemented based on the standard DiT [28], as shown in Fig.3(a). We introduce interleaved spatial and temporal attention to construct a spatio-temporal block, in which two new components, ACor and STM, are designed to enhance time series conditional guidance and explore complex spatio-temporal correlations. Details are provided in Sections III-D and III-E.
Embedding of Task-specific Conditions and Metadata. In UniTS, the task-specific conditions, target time series samples, and metadata corresponding to the four time series tasks are shown in Table III. We introduce a patch embedding layer for the input intermediate state ${\bf{x}}_{t}$} at flow time $t$ and condition ${\bf{x}}_{\text{con}}$}. Unlike previous latent transformers [29, 30, 31] that operate on latents encoded by variational autoencoder (VAE), UniTS directly tokenizes raw pixel inputs. We adapt the patch embedding from Vision Transformer (ViT) to each image in the time series, obtaining tokens ${\bf{z}}=\left\{{{{\bf{z}}_{t}},{{\bf{z}}_{\text{con}}}}\right\}\subset\mathbb{R}{{}^{T\!\times\!d\!\times\!{n_{\text{h}}}\!\times\!{n_{\text{w}}}}}$, where $d$ represents the channel dimension of each token, $n_{\text{h}}$} and $n_{\text{w}}$} are equal to $\frac{H}{h}$ and $\frac{W}{w}$, respectively. This process extracts non-overlapping image patches of size $h\times w$ from the time series for encoding. The tokens are then combined with 2D sincos positional encoding and fed into the spatio-temporal block.
Spatial and Temporal Positional Embeddings & Flow Time Embedding. Spatial and temporal positional embeddings are respectively added to the tokens in the first layer of the spatio-temporal block, enabling the model to understand the potential relationships of time series in both spatial and temporal dimensions. We introduce learnable spatial and temporal positional embeddings ${{\bf{p}}_{\text{spa}}}\in\mathbb{R}{{}^{{n_{\text{h}}}{n_{\text{w}}}\!\times d}},{{\bf{p}}_{\text{tmp}}}\in\mathbb{R}{{}^{T\!\times d}}$, which are initialized using 2D sincos positional encoding on a 2D grid of size ${n_{\text{h}}}\!\times\!{n_{\text{w}}}$, and 1D sincos positional encoding on a 1D time sequence of length $T$ respectively. After initialization, the spatial dimension of the spatial positional embedding is reshaped to ${n_{\text{h}}}{n_{\text{w}}}$. During training, a flow time $t$ is randomly sampled from the interval $[0,1]$ as the timestep and then transformed into sinusoidal timestep embedding ${{\bf{z}}_{\scriptscriptstyle\text{FM}}}\in\mathbb{R}{{}^{d}}$. To enable the spatio-temporal block to construct an adaptive velocity field along the continuous path defined by the flow time, the timestep embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}$} is repeated in both spatial and temporal dimensions to obtain ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}\in\mathbb{R}{{}^{{n_{\text{h}}}{n_{\text{w}}}\!\times\!d}}$ and ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{t}}\in\mathbb{R}{{}^{T\!\times d}}$, respectively.
Spatio-temporal Block. To enable UniTS to collaboratively model spatio-temporal dependencies, we introduce a spatio-temporal block with interleaved spatial and temporal attention, which captures geographic relationships in the spatial dimension and evolutionary patterns along the temporal dimension. The spatio-temporal block is structured sequentially from a spatial block to a temporal block, with each block consisting of ACor, adaptive layer normalization (AdaLN), and a spatial/temporal attention module incorporating STM, as illustrated in Fig.3(a). First, ${\bf{z}}_{t}$} and ${\bf{z}}_{\text{con}}$} are reshaped into $\left\{{{\bf{z}}_{t}^{\text{s}},{\bf{z}}_{\text{con}}^{\text{s}}}\right\}\subset\mathbb{R}{{}^{T\!\times\!{n_{\text{h}}}{n_{\text{w}}}\!\times\!d}}$ as input to the spatial block. In ACor, ${\bf{z}}_{\text{con}}^{\text{s}}$ is injected into ${\bf{z}}_{t}^{\text{s}}$} along the spatial dimension via condition feature-wise affine transformation, with the geographic embedding ${\bf{z}}_{\text{\tiny lonlat}}$ applied. Subsequently, the spatial flow time embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}$ is mapped through a linear layer to obtain scale $\gamma_{\text{s}}$, shift $\beta_{\text{s}}$, and gating parameter $\alpha_{\text{s}}$. The parameters $\gamma_{\text{s}}$ and $\beta_{\text{s}}$ are fed into AdaLN to adjust the feature flow in the model under the guidance of flow time. The adjusted features are passed into the spatial attention module, and then the output obtained by applying gating parameter is added to the input ${\bf{z}}_{t}^{\text{s}}$ through a residual connection to get ${\bf{\hat{z}}}_{t}^{\text{s}}$. The workflow of the spatial block is formulated as follows,
where $\rm{TMSA}(\cdot)$ represents temporal multi-head self-attention module. After the multi-layer spatio-temporal block, we obtain the generative time series clips ${\bf{\hat{x}}}_{t}$ at flow time $t$ by using a linear decoder with AdaLN and the unpatchify operation.
@@
Time Series
Semantic Change Detection
DynamicEarthNet Planet (RGBN), $C_{\text{con}}=4$ Segmentation map, $C=6$ -

### v2

# UniTS: Unified Spatio-Temporal Generative Model for Remote Sensing

@@ Abstract
One of the primary objectives of satellite remote sensing is to capture the complex dynamics of the Earth environment, which encompasses tasks such as reconstructing continuous cloud-free image sequences, detecting land cover changes, and forecasting future surface evolution. However, existing methods typically require specialized models tailored to different tasks, and lack a general framework that can address these multi-level tasks from a unified perspective. In this paper, we propose a Unified Spatio-Temporal Generative Model (UniTS), which integrates several long-separated core tasks, including time series reconstruction, time series cloud removal, time series semantic change detection, and time series forecasting. Based on the flow matching generative paradigm, UniTS constructs a deterministic evolution path from noise to targets under the guidance of task-specific conditions, achieving unified modeling of spatiotemporal representations for multi-level tasks. The UniTS architecture consists of a diffusion transformer with spatiotemporal blocks, where we design an Adaptive Condition Injector (ACor) to enhance the model’s conditional perception of multimodal inputs, enabling high-quality controllable generation. Additionally, we design a Spatiotemporal-aware Modulator (STM) to improve the ability of spatiotemporal blocks to capture complex spatiotemporal dependencies. Furthermore, we construct two high-quality multimodal time series datasets, TS-S12 and TS-S12CR, filling the gap of benchmark datasets for time series cloud removal and forecasting tasks. Extensive experiments demonstrate that UniTS exhibits exceptional generative and cognitive capabilities across reconstruction, interpretation, and forecasting tasks. It substantially outperforms existing specialized models, particularly under challenging conditions such as severe cloud contamination, modality absence, and forecasting complex phenological variations. More details can be found on the project page: https://yuxiangzhang-bit.github.io/UniTS-website/.
@@ I Introduction
Satellite image time series have become an indispensable tool for understanding and monitoring the dynamics of Earth’s systems, as they provide continuous spatiotemporal observational data. These long-term and consistent records of land surface dynamics deliver critical support for numerous fields, including ecological environment assessment, climate mitigation, and emergency response to natural disasters. With the increasing availability of high spatiotemporal resolution satellite data, time series analysis has expanded beyond single-sensor or single-modal approaches to integrate multispectral, radar, panchromatic, and other multi-source data, enabling more comprehensive and accurate perception of land surface changes.
@@
Low-level tasks such as gap filling and cloud removal are essential in optical remote sensing due to frequent atmospheric interference. Clouds, haze, and cloud shadows often block or distort optical signals, causing missing or corrupted data. According to statistics from long-term observations by the MODIS sensor [1], approximately 67% of the Earth’s surface and 55% of the land area are persistently affected by cloud cover on average. This high coverage rate not only significantly reduces the practical observation efficiency of optical satellites but also makes completely cloud-free time series images extremely scarce at regional scales. Additional issues like sensor failures or transmission errors can further degrade data quality. The primary goal of time series reconstruction is to recover complete, high-quality image time series from incomplete or noisy observational data. Common approaches use cloud masks to simulate data gaps on clean sequences. Stucker et al. [2] introduced U-TILISE, a representation learning framework using an encoder-decoder with temporal attention. From a generation viewpoint, Shu et al. [3] developed RESTORE-Diffusion Transformer (DiT), a multimodal diffusion framework that fuses optical and Synthetic Aperture Radar (SAR) time series with date embeddings to guide reconstruction under clouds.
@@
Time series cloud removal is closely related to time series reconstruction and poses greater challenges. Unlike simply reconstructing areas obscured by cloud masks, this task specifically aims at restoring pixels occluded by clouds and their shadows within a temporal sequence. Patrick et al. [4] constructed the SEN12MS-CR-TS dataset, providing 30 temporally aligned Sentinel-1/2 image pairs with corresponding cloud masks. To address data gaps in visible and near-infrared bands, Gonzalez-Calabuig et al. [5] proposed GANFilling, a GAN-based method that integrates convolutional LSTM layers to convert cloudy optical sequences into cloud-free ones.
@@
The continuous spatiotemporal time series data used for understanding and analysis in high-level tasks is provided by low-level tasks. Time series semantic change detection monitors complex land cover dynamics by identifying not only the precise location (Where) and timing (When) of changes but also provide semantic information about the type of change (What). Garnot et al. [6] proposed a U-Net with Temporal Attention Encoder (U-TAE) for time series panoptic segmentation. However, this method is only designed to predict a single land cover map corresponding to the time series images and does not support land cover change detection. He et al. [7] introduced an end-to-end framework based on a fully convolutional network, which directly learns the mapping between spectral features and land cover classes to achieve time series semantic change detection in a unified manner.
Time series forecasting predicts future surface conditions by analyzing historical Earth observation data and related factors, requiring models to capture temporal evolution patterns, seasonal trends, and long-term spatiotemporal dependencies. One early benchmark is the EarthNet2021 dataset by Requena‑Mesa et al. [8], which pairs Sentinel‑2 imagery with topographic and meteorological variables for surface forecasting. Currently, Benson et al. [9] introduced the GreenEarthNet dataset for high-resolution vegetation forecasting and developed a multimodal transformer model named Contextformer.
@@
  • Limited exploration in time series cloud removal tasks and challenges in dataset construction. Most studies focus on reconstructing time series using cloud masks, which oversimplify real-world conditions with variable cloud contamination. Progress in time series cloud removal is limited, and the currently available benchmark datasets exhibit significant shortcomings. For example, the SEN12MS-CR-TS dataset [4] suffers from temporal misalignment of up to 14 days and excludes images with more than 50% cloud cover, while the EarthNet2021 dataset [8] also omits heavily clouded sequences. Consequently, the absence of high-quality, temporally aligned time series datasets severely hinders reliable model training and evaluation in this field.
@@
  • Limited exploration in time series forecasting tasks. Most existing studies employ discriminative models, such as ConvLSTM and 3D CNN, exemplified by Pangu-Weather [10]. However, these models struggle to capture the complexities of spatiotemporal data. There is a need to explore generative models in time series forecasting research that can better fit complex distributions. Current studies mainly focus on predicting vegetation indices, with limited work on forecasting original remote sensing images, especially for high-resolution multispectral data like Sentinel-2. Due to its high spectral dimensionality and complex spatial details, predicting original multispectral time series presents even greater challenges.
@@
  • Lack of a unified framework for multiple remote sensing time series tasks. Current research remains largely in the stage of developing specialized models tailored for specific tasks, and a unified framework that can effectively address diverse time series tasks is still missing. We argue that the core challenge for both low-level and high-level time series tasks lies in mining spatiotemporal representations, with the main difference being how these representations are adapted through task-specific designs to meet particular requirements.
@@
To address the aforementioned technical and data gaps, we propose Unified Spatio-Temporal Generative Model (UniTS) based on the flow matching-based generative paradigm, which is applicable to four time series tasks, as shown in Fig.1. Additionally, we construct two high-quality benchmarks TS-S12 and TS-S12CR for time series model validation. The main contributions of this paper are as follows:
@@
  • UniTS achieves a unified modeling for the first time in multi-level tasks: time series reconstruction, time series cloud removal, time series semantic change detection, and time series forecasting. The framework demonstrates excellent generation and understanding capabilities across different task levels, whether it is low-level time series reflectance recovery or high-level time series semantic understanding.
@@
  • Within the diffusion transformer framework, we introduce a spatiotemporal block and design two novel components: the Adaptive Condition Injector (ACor) and the Spatiotemporal-aware Modulator (STM). ACor adaptively injects multimodal conditional information (e.g., SAR and optical imagery) by dynamically generating affine transformation parameters, significantly enhancing the model’s conditional perception of multimodal inputs across various time series tasks. Meanwhile, STM modulates attention weights in the spatiotemporal block by leveraging generated dynamic bias terms based on spatiotemporal priors, thereby strengthening the model’s capacity to capture complex spatiotemporal dependencies.
@@
  • We construct two high-quality multimodal time series datasets, namely TS-S12 and TS-S12CR. Among them, TS-S12 and TS-S12CR contain Sentinel-1 and Sentinel-2 imagery from 14,973 and 12,126 ROIs around the world, respectively. TS-S12 provides aligned sample pairs of Sentinel-1 and cloud-free Sentinel-2 for time series reconstruction and forecasting tasks. TS-S12CR offers aligned triplet samples of Sentinel-1, cloud-covered Sentinel-2, and cloud-free Sentinel-2 specifically designed for time series cloud removal task. TS-S12CR provides an extreme scenario with an average cloud coverage of 84.02%, serving as an important benchmark for developing robust time series cloud removal methods.
@@ III Unified Time Series Generative Model
In this section, we first introduce the framework of UniTS in Section III-C. The details of ACor and STM are presented in Section III-D and III-E, respectively. Finally, we elaborate on the training and inference processes for conditional time series generation in Section III-F. Notations used in this paper are summarized in Table II.
@@ III-C UniTS Framework
Time Series Semantic Change Detection DynamicEarthNet Planet (RGBN), $C_{\text{con}}=4$ Segmentation map, $C=6$ -
@@
Architecture. We instantiate ${f_{\theta}}(\cdot)$ in Eq.(2) using a transformer-based architecture. This architecture is chosen for its simplicity, scalability, and effectiveness in generative modeling. Specifically, UniTS is implemented based on the standard DiT [21], as shown in Fig.3(a). We introduce interleaved spatial and temporal attention to construct a spatiotemporal block, in which two new components, ACor and STM, are designed to enhance time series conditional guidance and explore complex spatiotemporal correlations. Details are provided in Sections III-D and III-E.
Embedding of Task-specific Conditions and Metadata. In UniTS, the task-specific conditions, target time series samples, and metadata corresponding to the four time series tasks are shown in Table III. We introduce a patch embedding layer for the input intermediate state ${\bf{x}}_{t}$} at flow time $t$ and condition ${\bf{x}}_{\text{con}}$}. Unlike previous latent transformers [22, 23, 24] that operate on latents encoded by variational autoencoder (VAE), UniTS directly tokenizes raw pixel inputs. We adapt the patch embedding from Vision Transformer (ViT) to each image in the time series, obtaining tokens ${\bf{z}}=\left\{{{{\bf{z}}_{t}},{{\bf{z}}_{\text{con}}}}\right\}\subset\mathbb{R}{{}^{T\!\times\!d\!\times\!{n_{\text{h}}}\!\times\!{n_{\text{w}}}}}$, where $d$ represents the channel dimension of each token, $n_{\text{h}}$} and $n_{\text{w}}$} are equal to $\frac{H}{h}$ and $\frac{W}{w}$, respectively. This process extracts non-overlapping image patches of size $h\times w$ from the time series for encoding. The tokens are then combined with 2D sincos positional encoding and fed into the spatiotemporal block.
Spatial and Temporal Positional Embeddings & Flow Time Embedding. Spatial and temporal positional embeddings are respectively added to the tokens in the first layer of the spatiotemporal block, enabling the model to understand the potential relationships of time series in both spatial and temporal dimensions. We introduce learnable spatial and temporal positional embeddings ${{\bf{p}}_{\text{spa}}}\in\mathbb{R}{{}^{{n_{\text{h}}}{n_{\text{w}}}\!\times d}},{{\bf{p}}_{\text{tmp}}}\in\mathbb{R}{{}^{T\!\times d}}$, which are initialized using 2D sincos positional encoding on a 2D grid of size ${n_{\text{h}}}\!\times\!{n_{\text{w}}}$, and 1D sincos positional encoding on a 1D time sequence of length $T$ respectively. After initialization, the spatial dimension of the spatial positional embedding is reshaped to ${n_{\text{h}}}{n_{\text{w}}}$. During training, a flow time $t$ is randomly sampled from the interval $[0,1]$ as the timestep and then transformed into sinusoidal timestep embedding ${{\bf{z}}_{\scriptscriptstyle\text{FM}}}\in\mathbb{R}{{}^{d}}$. To enable the spatiotemporal block to construct an adaptive velocity field along the continuous path defined by the flow time, the timestep embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}$} is repeated in both spatial and temporal dimensions to obtain ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}\in\mathbb{R}{{}^{{n_{\text{h}}}{n_{\text{w}}}\!\times\!d}}$ and ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{t}}\in\mathbb{R}{{}^{T\!\times d}}$, respectively.
spatiotemporal Block. To enable UniTS to collaboratively model spatiotemporal dependencies, we introduce a spatiotemporal block with interleaved spatial and temporal attention, which captures geographic relationships in the spatial dimension and evolutionary patterns along the temporal dimension. The spatiotemporal block is structured sequentially from a spatial block to a temporal block, with each block consisting of ACor, adaptive layer normalization (AdaLN), and a spatial/temporal attention module incorporating STM, as illustrated in Fig.3(a). First, ${\bf{z}}_{t}$} and ${\bf{z}}_{\text{con}}$} are reshaped into $\left\{{{\bf{z}}_{t}^{\text{s}},{\bf{z}}_{\text{con}}^{\text{s}}}\right\}\subset\mathbb{R}{{}^{T\!\times\!{n_{\text{h}}}{n_{\text{w}}}\!\times\!d}}$ as input to the spatial block. In ACor, ${\bf{z}}_{\text{con}}^{\text{s}}$ is injected into ${\bf{z}}_{t}^{\text{s}}$} along the spatial dimension via condition feature-wise affine transformation, with the geographic embedding ${\bf{z}}_{\text{\tiny lonlat}}$ applied. Subsequently, the spatial flow time embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}$ is mapped through a linear layer to obtain scale $\gamma_{\text{s}}$, shift $\beta_{\text{s}}$, and gating parameter $\alpha_{\text{s}}$. The parameters $\gamma_{\text{s}}$ and $\beta_{\text{s}}$ are fed into AdaLN to adjust the feature flow in the model under the guidance of flow time. The adjusted features are passed into the spatial attention module, and then the output obtained by applying gating parameter is added to the input ${\bf{z}}_{t}^{\text{s}}$ through a residual connection to get ${\bf{\hat{z}}}_{t}^{\text{s}}$. The workflow of the spatial block is formulated as follows,
where $\rm{TMSA}(\cdot)$ represents temporal multi-head self-attention module. After the multi-layer spatiotemporal block, we obtain the generative time series clips ${\bf{\hat{x}}}_{t}$ at flow time $t$ by using a linear decoder with AdaLN and the unpatchify operation.